### v1

@@ 3 Method
The input to the proposed method is a video sample of a student seated in front of a laptop or PC camera during a virtual learning session. The sequences of facial landmarks extracted from consecutive video frames through MediaPipe [lugaresi2019mediapipe, grishchenko2020attention] are analyzed by ST-GCN [yan2018spatial, 10144079] to output the engagement level of the student in the video.
@@ 3.1 Graph-based Representation
MediaPipe [lugaresi2019mediapipe], an open-source framework developed by Google, facilitates the construction of applied machine-learning pipelines for processing video data. The framework’s capabilities extend to a variety of tasks such as face tracking, hand and pose estimation, and facial landmark estimation, offering pre-trained models and tools suitable for real-time and streaming applications [latreche2023reliability]. Grishchenko et al. [grishchenko2020attention] introduced Attention Mesh, which was incorporated into MediaPipe [lugaresi2019mediapipe]. Attention Mesh is a pre-trained, end-to-end 3D facial landmark detection model that is both efficient and lightweight. It is enhanced with an attention mechanism focused on areas of semantic significance, such as the eyes, iris, and lips. Attention Mesh is capable of identifying 468 3D facial landmarks across the entire face and an additional 10 landmarks specifically for the iris. However, for the proposed method for engagement measurement, not all 478 landmarks are utilized. In line with existing research [liu2022graph], 68 3D facial landmarks that correspond to those identified by Dlib’s [king_dlib] framework, along with 10 3D landmarks from the iris, totaling 78 landmarks are utilized.
The 3D facial landmarks encapsulate crucial spatial-temporal information pertinent to head pose [malek2021head], AUs [jacob2021facial], eye gaze [grishchenko2020attention], and affect [liu2022graph, toisoul2021estimation]—key features identified in prior engagement measurement studies [karimah2022automatic, dewan2019engagement, khan2022inconsistencies, chen2019faceengage, wu2020advanced, ma2021automatic, copur2022engagement, abedi2023affect, abedi2023detecting, thomas2018predicting, thomas2022automatic, gupta2016daisee, zhang2019novel, abedi2021improving, mehta2022three]. Consequently, there is no necessity for extracting additional handcrafted features from the video frames.
The $N$ 3D facial landmarks extracted from $T$ consecutive video frames are utilized to construct a spatiotemporal graph $G=(V,E)$. In this graph, the set of nodes $V=\{v_{ti}|t=1,\ldots,T,i=1,\ldots,N\}$ encompasses all the facial landmarks in a sequence. To construct $G$, first, the facial landmarks within one frame are connected with edges according to a connectivity structure based on Delaunay triangulation [liu2019facial] which is consistent with true facial muscle distribution and uniform for different subjects [liu2022graph]. Then each landmark will be connected to the same landmark in the consecutive frame.
@@ 3.2 Graph-based Reasoning
Based on the spatiotemporal graph of facial landmarks $G=(V,E)$ constructed above, an adjacency matrix $A$ is defined as an $N\times N$ matrix where the element at position $(i,j)$ is set to 1 if there is an edge connecting the $i^{th}$ and $j^{th}$ landmarks, and set to 0 otherwise. An identity matrix $I$, of the same dimensions as $A$, is created to represent self-connections. A spatial-temporal graph, as the basic element of an ST-GCN layer, is implemented as follows [kipf2016semi, yan2018spatial].
@@
where $\Lambda^{ii}=\sum_{j}(A^{ij}+I^{ij})$. $M$ is a learnable weight matrix that enables scaling the contributions of a node’s feature to its neighboring nodes [yan2018spatial]. The input feature map, denoted $f_{\text{in}}$, is the raw coordinates of facial landmarks for the first layer of ST-GCN, and it represents the outputs from previous layers in subsequent layers of ST-GCN. The dimensionality of $f_{\text{in}}$ is $(C,N,T)$, where $C$ is the number of channels, for example, 3 in the initial input to the network corresponding to the $x$, $y$, and $z$ coordinates of the facial landmarks. In each layer of ST-GCN, initially, the spatial (intra-frame) convolution is applied to $f_{\text{in}}$ based on the weight matrix $W_{\text{spatial}}$, utilizing a standard 2D convolution with a kernel size of $1\times 1$. Subsequently, the resulting tensor is multiplied by the normalized adjacency matrix $\Lambda^{-\frac{1}{2}}((A+I)\odot M)\Lambda^{-\frac{1}{2}}$ across the spatial dimension. Afterward, the temporal (inter-frame) convolution, based on the weight matrix $W_{\text{temporal}}$, is applied to the tensor output from the spatial convolution. This convolution is a standard 2D convolution with a kernel size of $1\times\Gamma$, where $\Gamma$ signifies the temporal kernel size.
Following an adequate number of ST-GCN layers that perform spatial-temporal graph convolutions as outlined above, the resulting tensor undergoes 2D average pooling, and the final output of the network is generated by a final 2D convolution. This convolution employs a kernel size of $1\times 1$ and features an output channel dimensionality equal to the number of classes $K$, i.e., the number of engagement levels to be measured.
@@ 3.3 Ordinal Engagement classification through Transfer Learning
The model described above, with a final layer comprising $K$ output channels, tackles the engagement measurement problem as a categorical $K$-class classification problem without taking into account the ordinal nature of the engagement variable [yannakakis2018ordinal, whitehill2014faces, khan2022inconsistencies]. The model could harness the ordinality of the engagement variable to enhance its inferences. Drawing inspiration from [frank2001simple, abedi2023affect], a novel ordinal learning framework based on transfer learning is introduced as follows.
@@
Inference phase– For the ordinal classification of a test sample, the sample is initially input into the pre-trained (and frozen) sequence of ST-GCN layers, followed by a 2D average pooling layer. The tensor obtained from this process is then input into $K-1$ pre-trained final 2D convolution layers, each yielding a probability estimate for the test sample being in the binary class $y_{t}=y_{i}$, where $i=0,1,\ldots,K-2$. Subsequently, these $K-1$ binary probability estimates are transformed into a single multi-class probability of the sample belonging to class $y=0,1,\ldots,K-1$, as follows [frank2001simple].
@@ 4 Experiments
This section evaluates the performance of the proposed method relative to existing methods in video-based engagement measurement. It reports and discusses the results of multi-class and binary classification of engagement across two datasets.
4.1 Evaluation Metrics
Based on the engagement measurement problem at hand, several evaluation metrics are employed. In the context of multi-class engagement level classification, metrics such as accuracy and confusion matrix are reported. For binary engagement classification, accuracy, Area Under the Curve of the Receiver Operating Characteristic (AUC-ROC), and Area Under the Curve of the Precision and Recall curve (AUC-PR) are utilized.
Table 1: The distribution of samples in not engaged, barely engaged, engaged, and highly engaged classes in the train and validation sets in the EngageNet dataset [engagenet].
Engagement Level Train Validation
Not engaged 1550 132
Barely engaged 1035 97
Engaged 1658 273
Highly Engaged 3740 569
Total 7983 1071
Table 2: The distribution of samples in engaged and not engaged classes in the train, validation, and test sets in the IIITB Online SE dataset [thomas2022automatic].
Engagement Level Train Validation Test
Not engaged 570 580 570
Engaged 2620 1080 720
Total 3190 1660 1290
@@ 4.2 Datasets
Experiments on two video-based engagement measurement datasets were conducted, each presenting unique challenges that further enabled the validation of the proposed method.
EngageNet The EngageNet dataset [engagenet], recognized as the largest publicly available dataset for video-based student engagement measurement, includes video recordings of 127 subjects participating in virtual learning courses. Each video sample has a duration of 10 seconds, with a frame rate of 30 fps and a resolution of $1280\times 720$ pixels. Three independent annotators annotated the subjects’ video recordings as four ordinal levels of engagement: not-engaged, barely engaged, engaged, and highly engaged. The inter-rater reliability among the annotators was high, with values of 0.73, 0.77, and 0.79, indicating a substantial agreement level. The dataset was divided into 7983, 1071, and 2257 samples for training, validation, and testing, respectively, using a subject-independent data split approach [engagenet]. However, only the training and validation sets were made available by the dataset creators and were utilized for the training and validation of predictive models in the experiments presented in this paper. Table 1 details the distribution of the four engagement classes within the training and validation sets.
Online SE: The Online SE dataset [thomas2022automatic], comprises videos of six students participating in online courses via the Zoom platform. These recordings span 10 seconds each, with a frame rate of 24 fps and a resolution of $220\times 155$ pixels. Two independent annotators annotated the video recordings as engaged and not engaged using the Human Expert Labeling Process (HELP) [aslan2017human, abedi2023bag]. The binary engagement labels achieved an inter-rater reliability score of 0.62, denoting a significant level of agreement. Following the exclusion of flawed videos [abedi2023bag], the dataset was segmented into 3190, 1660, and 1290 samples for training, validation, and testing, respectively. The distribution of engaged and not-engaged samples across the training, validation, and test sets is outlined in Table 2.
@@ 4.3 Experimental Setting
The sole information extracted from video frames is facial landmarks, which are analyzed by ST-GCN to determine the engagement level of the student in the video. Drawing inspiration from the pioneering works on body-joints-based action analysis [zheng2023skeleton, yan2018spatial], the proposed ST-GCN for facial-landmarks-based engagement measurement is structured as follows. The input facial landmarks are first processed through a batch normalization layer, followed by three consecutive ST-GCN layers with 64, 128, and 256 output channels, respectively. Residual connections are incorporated in the last two ST-GCN layers. A dropout rate of 0.1 is applied to each ST-GCN layer. The temporal kernel size in the ST-GCN layers is selected to be $9$. Subsequent to the ST-GCN layers, an average pooling layer is utilized, and its resulting tensor is directed into a 2D convolutional layer with 256 input channels and a number of output channels corresponding to the number of classes. A Softmax activation function then computes the probability estimates. In cases where engagement measurement is framed as a binary classification task, the terminal 2D convolution layer is configured with a single output channel, substituting Softmax with a Sigmoid function. The Sigmoid function is also employed for the individual binary classifiers within the ordinal learning framework detailed in subsection 3.3. The models are trained using the Adam optimizer with mini-batches of size $16$ and a learning rate of $0.001$ for $300$ epochs. The learning rate is decayed by a factor of $0.1$ every $100$ epochs.
Table 3: Classification accuracy of engagement levels on the validation set of the EngageNet dataset [engagenet]: comparison of various feature sets and classification models from [engagenet] against two configurations of the proposed method - facial landmarks analyzed by ST-GCN and ordinal ST-GCN. Bolded values denote the best results.
Features Model Accuracy
Gaze LSTM 0.6125
Head Pose LSTM 0.6760
AU LSTM 0.6303
Gaze + Head Pose LSTM 0.6769
Gaze + Head Pose + AU LSTM 0.6704
Gaze CNN-LSTM 0.6060
Head Pose CNN-LSTM 0.6732
AU CNN-LSTM 0.6172
Gaze + AU CNN-LSTM 0.6275
Head Pose + AU CNN-LSTM 0.6751
Gaze + Head Pose + AU CNN-LSTM 0.6751
Gaze TCN 0.6256
Head Pose TCN 0.6611
AU TCN 0.6293
Gaze + Head Pose + AU TCN 0.6779
Gaze Transformer 0.5545
Gaze + Head Pose Transformer 0.6445
Gaze + Head Pose + AU Transformer 0.6910
Gaze + Head Pose + AU + MARLIN Transformer 0.6849
Facial Landmarks ST-GCN 0.6937
Facial Landmarks Ordinal ST-GCN 0.7124
the proposed method
@@ 4.4 Experimental Results
Table 3 presents the comparative results of two settings of the proposed method, a regular non-ordinal classifier and an ordinal classifier, with previous methods on the validation set of the EngageNet dataset [engagenet]. The engagement measurement in EngageNet [engagenet] is a four-class classification problem and the accuracy is reported as the evaluation metric. The previous methods used different combinations of OpenFace’s eye gaze, head pose, and AU features [baltrusaitis2018openface] along with MARLIN’s facial embedding features [cai2023marlin]. These features were classified by LSTM, CNN-LSTM, TCN, and Transformer, refer to [engagenet] for more details of the previous methods.
@@
In the single feature configurations of previous methods in Table 3, head pose achieves better results compared to AUs, which is better than eye gaze. While head pose and eye gaze are indicators of behavioral engagement [abedi2023affect, ocumpaugh2015baker], AUs, which are associated with facial expressions and affect [toisoul2021estimation], are indicators of affective engagement. Combining these three features is always beneficial since engagement is a multi-component variable that can be measured by affective and behavioral indicators when the only available data is video [abedi2023affect]. Among classification models, utilizing more advanced models improves accuracy; the Transformer is better than TCN, which is better than CNN-LSTM and LSTM; however, it comes at the cost of increased computational complexity and the number of model parameters. It should also be noted that the feature extraction step for the three feature sets requires three complex deep learning models in OpenFace [baltrusaitis2018openface]. Considering only the classification models, the number of parameters in the Transformer, according to the architecture described in [engagenet], is 104,963,076, while the numbers of parameters in the proposed non-ordinal and ordinal ST-GCN are 861,688, and 861,431, respectively, indicating the efficiency of the proposed method. While being lightweight, the proposed method with ordinal STGCN improves the state-of-the-art by 3.1% compared to eye gaze, head pose, and AUs with Transformer.
@@
Tables 4(a) and 4(b) depict the confusion matrices of the two configurations of the proposed method in Table 3, facial landmarks classified by non-ordinal and ordinal ST-GCN, respectively. As shown in Tables 4(a) and 4(b), incorporating ordinality significantly increases the number of correctly classified samples in the first three classes and results in a 2.7% improvement in accuracy compared to its non-ordinal counterpart.
@@
Table 4: Confusion matrices of the proposed method using facial landmarks with (a) non-ordinal and (b) ordinal ST-GCN on the validation set of the EngageNet dataset [engagenet].
@@
Table 5 displays the results of different variants of the proposed method on the validation set of the EngageNet [engagenet] dataset. In the first two variants, the $x$, $y$, and $z$ coordinates of facial landmarks are converted into multivariate time series and analyzed by an LSTM and TCN. The LSTM includes four unidirectional layers with 256 neurons in hidden units and is followed by a $256\times 4$ fully connected layer. The parameters of the TCN are as follows: the number of layers, number of filters, kernel size, and dropout rate are 8, 64, 8, and 0.05, respectively. While their results are acceptable and better than most of the earlier methods in Table 3, they cannot outperform ST-GCN, the last two columns of Table 3. The fact that the accuracy of the LSTM and TCN in Table 5 is higher than those in Table 3 signifies the efficiency of facial landmarks for engagement measurement. In the third variant, the $z$ coordinates of facial landmarks are disregarded, and the decrease in accuracy indicates the importance of the $z$ coordinates for engagement measurement. Temporal kernel sizes other than 9 in the fourth and fifth variants have a negative impact on the results. When engagement measurement is performed using every 2, 4, 8, and 16 frames, instead of every frame, there is a slight decrease in accuracy. However, this is a trade-off between accuracy and computation since reducing the frame rate corresponds to a reduction in computation.
Table 5: Classification accuracy of engagement levels on the validation set of the EngageNet dataset [engagenet] for different variants of the proposed method.
@@
Table 6 presents the results of the proposed method in comparison to previous methods on the test set of the Online SE dataset [thomas2022automatic]. The earlier methods listed in Table 6 are feature-based, extracting affective and behavioral features from video frames and performing binary classification of the features using TCN in [thomas2018predicting], TCN in [abedi2023affect], LSTM with attention in [chen2019faceengage], and BoW in [abedi2023bag]. Given that engagement measurement in the Online SE dataset [thomas2022automatic] is posed as a binary classification problem, implementing the ordinal version of the proposed method is not required. The proposed method, employing facial landmarks with ST-GCN, attains the highest accuracy and AUC PR.
Table 6: Binary classification accuracy, AUC ROC, and AUC PR of engagement on the test set of the Online SE dataset [thomas2022automatic]: comparison of previous methods against the proposed method. Bolded values denote the best results.

### v2

@@ 3 Method
The input to the proposed method is a video sample of a student seated in front of a laptop or PC camera during a virtual learning session. The sequences of facial landmarks extracted from consecutive video frames through MediaPipe [14, 20] are analyzed by ST-GCN [21, 16] to output the engagement level of the student in the video.
@@ 3.1 Graph-based Representation
The MediaPipe deep learning solution [14], a framework that is both real-time and cross-platform, is employed for the extraction of facial landmarks from video. Incorporated within MediaPipe, Attention Mesh [20] is adept at detecting 468 3D facial landmarks throughout the face and an extra 10 landmarks for the iris. However, not all 478 landmarks are employed in the proposed engagement measurement method. Consistent with existing studies [3], only 68 of the 3D facial landmarks, which match those identified by the Dlib framework [37], in addition to the 10 3D iris landmarks, making a total of 78 landmarks, are utilized.
The 3D facial landmarks encapsulate crucial spatial-temporal information pertinent to head pose [17], AUs [18], eye gaze [20], and affect [3, 19]—key features identified in prior engagement measurement studies [8, 9, 10, 22, 23, 13, 24, 33, 26, 27]. Consequently, there is no necessity for extracting additional handcrafted features from the video frames.
The $N$ 3D facial landmarks extracted from $T$ consecutive video frames are utilized to construct a spatiotemporal graph $G=(V,E)$. In this graph, the set of nodes $V=\{v_{ti}|t=1,\ldots,T,i=1,\ldots,N\}$ encompasses all the facial landmarks in a sequence. To construct $G$, first, the facial landmarks within one frame are connected with edges according to a connectivity structure based on Delaunay triangulation [38] which is consistent with true facial muscle distribution and uniform for different subjects [3]. Then each landmark will be connected to the same landmark in the consecutive frame.
@@ 3.2 Graph-based Reasoning
Based on the spatiotemporal graph of facial landmarks $G=(V,E)$ constructed above, an adjacency matrix $A$ is defined as an $N\times N$ matrix where the element at position $(i,j)$ is set to 1 if there is an edge connecting the $i^{th}$ and $j^{th}$ landmarks, and set to 0 otherwise. An identity matrix $I$, of the same dimensions as $A$, is created to represent self-connections. A spatial-temporal graph, as the basic element of an ST-GCN layer, is implemented as follows [21].
@@
where $\Lambda^{ii}=\sum_{j}(A^{ij}+I^{ij})$. $M$ is a learnable weight matrix that enables scaling the contributions of a node’s feature to its neighboring nodes [21]. The input feature map, denoted $f_{\text{in}}$, is the raw coordinates of facial landmarks for the first layer of ST-GCN, and it represents the outputs from previous layers in subsequent layers of ST-GCN. The dimensionality of $f_{\text{in}}$ is $(C,N,T)$, where $C$ is the number of channels, for example, 3 in the initial input to the network corresponding to the $x$, $y$, and $z$ coordinates of the facial landmarks. In each layer of ST-GCN, initially, the spatial (intra-frame) convolution is applied to $f_{\text{in}}$ based on the weight matrix $W_{\text{spatial}}$, utilizing a standard 2D convolution with a kernel size of $1\times 1$. Subsequently, the resulting tensor is multiplied by the normalized adjacency matrix $\Lambda^{-\frac{1}{2}}((A+I)\odot M)\Lambda^{-\frac{1}{2}}$ across the spatial dimension. Afterward, the temporal (inter-frame) convolution, based on the weight matrix $W_{\text{temporal}}$, is applied to the tensor output from the spatial convolution. This convolution is a standard 2D convolution with a kernel size of $1\times\Gamma$, where $\Gamma$ signifies the temporal kernel size.
Following an adequate number of ST-GCN layers, specifically three in this work, that perform spatial-temporal graph convolutions as outlined above, the resulting tensor undergoes 2D average pooling. The final output of the network is generated by a final 2D convolution. This convolution employs a kernel size of $1\times 1$ and features an output channel dimensionality equal to the number of classes $K$, i.e., the number of engagement levels to be measured. An explanation of the detailed architecture of the ST-GCNs for specific datasets can be found in subsection 4.2.
@@ 3.3 Ordinal Engagement classification through Transfer Learning
The model described above, with a final layer comprising $K$ output channels, tackles the engagement measurement problem as a categorical $K$-class classification problem without taking into account the ordinal nature of the engagement variable [39, 40, 10]. The model could harness the ordinality of the engagement variable to enhance its inferences. Drawing inspiration from [41, 13], a novel ordinal learning framework based on transfer learning is introduced as follows.
@@
Inference phase– For the ordinal classification of a test sample, the sample is initially input into the pre-trained (and frozen) sequence of ST-GCN layers, followed by a 2D average pooling layer. The tensor obtained from this process is then input into $K-1$ pre-trained final 2D convolution layers, each yielding a probability estimate for the test sample being in the binary class $y_{t}=y_{i}$, where $i=0,1,\ldots,K-2$. Subsequently, these $K-1$ binary probability estimates are transformed into a single multi-class probability of the sample belonging to class $y=0,1,\ldots,K-1$, as follows [41].
@@ 4 Experiments
This section evaluates the performance of the proposed method relative to existing methods in video-based engagement measurement. It reports and discusses the results of multi-class and binary classification of engagement across two datasets. Based on the engagement measurement problem at hand, several evaluation metrics are employed. In the context of multi-class engagement level classification, metrics such as accuracy and confusion matrix are reported. For binary engagement classification, accuracy, the Area Under the Curve of the Receiver Operating Characteristic (AUC-ROC), and the Area Under the Curve of the Precision and Recall curve (AUC-PR) are utilized. In addition, the number of parameters of the models used, memory consumption, and inference time of the proposed method are compared to those of previous methods.
@@ 4.1 Datasets
Experiments on two large video-based engagement measurement datasets were conducted, each presenting unique challenges that further enabled the validation of the proposed method.
EngageNet: The EngageNet dataset [12], recognized as the largest dataset for student engagement measurement, includes video recordings of 127 subjects participating in virtual learning sessions. Each video sample has a duration of 10 seconds, with a frame rate of 30 fps and a resolution of $1280\times 720$ pixels. The subjects’ video recordings were annotated as four ordinal levels of engagement: Not-Engaged, Barely-Engaged, Engaged, and Highly-Engaged. The dataset was divided into 7983, 1071, and 2257 samples for training, validation, and testing, respectively, using a subject-independent data split approach [12]. However, only the training and validation sets were made available by the dataset creators and were utilized for the training and validation of predictive models in the experiments presented in this paper. The distribution of samples in the four aforementioned classes of engagement in the training and validation sets are 1550, 1035, 1658, and 3740, and 132, 97, 273, and 569, respectively.
Online SE: The Online SE dataset [42], comprises videos of six students participating in online courses via the Zoom platform. These recordings span 10 seconds each, with a frame rate of 24 fps and a resolution of $220\times 155$ pixels. The videos were annotated as either Not-Engaged or Engaged. The dataset was segmented into 3190, 1660, and 1290 samples for training, validation, and testing, respectively. The distribution of samples in the two aforementioned classes of engagement in the training, validation, and test sets is 570 and 2620, 580 and 1080, and 570 and 720, respectively.
@@ 4.2 Experimental Setting
The sole information extracted from video frames is facial landmarks, which are analyzed by ST-GCN to determine the engagement level of the student in the video. Drawing inspiration from the pioneering works on body-joints-based action analysis [21, 43], the proposed ST-GCN for facial-landmarks-based engagement measurement is structured as follows. The input facial landmarks are first processed through a batch normalization layer, followed by three consecutive ST-GCN layers with 64, 128, and 256 output channels, respectively. Residual connections are incorporated in the last two ST-GCN layers. A dropout rate of 0.1 is applied to each ST-GCN layer. The temporal kernel size in the ST-GCN layers is selected to be $9$. Subsequent to the ST-GCN layers, an average pooling layer is utilized, and its resulting tensor is directed into a 2D convolutional layer with 256 input channels and a number of output channels corresponding to the number of classes. A Softmax activation function then computes the probability estimates. In cases where engagement measurement is framed as a binary classification task, the terminal 2D convolution layer is configured with a single output channel, substituting Softmax with a Sigmoid function. The Sigmoid function is also employed for the individual binary classifiers within the ordinal learning framework detailed in subsection 3.3. The models are trained using the Adam optimizer with mini-batches of size $16$ and an initial learning rate of $0.001$ for $300$ epochs. The learning rate is decayed by a factor of $0.1$ every $100$ epochs.
Table 1: Classification accuracy of engagement levels on the validation set of the EngageNet dataset [12]: comparison of state-of-the-art end-to-end methods and feature-based methods with various feature sets and classification models against two configurations of the proposed method - facial landmarks analyzed by ST-GCN and ordinal ST-GCN. Bolded values denote the best results.
Ref. Features Model Accuracy
[27] End to End Model ResNet + TCN 0.5472
[30] End to End Model EfficientNet + LSTM 0.5757
[30] End to End Model EfficientNet + Bi-LSTM 0.5894
[12] Gaze LSTM 0.6125
[12] Head Pose LSTM 0.6760
[12] AU LSTM 0.6303
[12] Gaze + Head Pose LSTM 0.6769
[12] Gaze + Head Pose + AU LSTM 0.6704
[12] Gaze CNN-LSTM 0.6060
[12] Head Pose CNN-LSTM 0.6732
[12] AU CNN-LSTM 0.6172
[12] Gaze + AU CNN-LSTM 0.6275
[12] Head Pose + AU CNN-LSTM 0.6751
[12] Gaze + Head Pose + AU CNN-LSTM 0.6751
[12] Gaze TCN 0.6256
[12] Head Pose TCN 0.6611
[12] AU TCN 0.6293
[12] Gaze + Head Pose + AU TCN 0.6779
[12] Gaze Transformer 0.5545
[12] Gaze + Head Pose Transformer 0.6445
[12] Gaze + Head Pose + AU Transformer 0.6910
[12] Gaze + Head Pose + AU + MARLIN Transformer 0.6849
[34] Gaze TCCT-Net 0.6433
[34] Head Pose TCCT-Net 0.6891
[34] AU TCCT-Net 0.6629
[34] Gaze + Head Pose TCCT-Net 0.6564
[34] Gaze + Head Pose + AU TCCT-Net 0.6713
Ours Facial Landmarks ST-GCN 0.6937
Ours Facial Landmarks Ordinal ST-GCN 0.7124
@@ 4.3.1 Comparison to Previous Methods
Table 1 presents the comparative results of two settings of the proposed method, a regular non-ordinal classifier and an ordinal classifier, with previous methods on the validation set of the EngageNet dataset [12]. The engagement measurement in EngageNet [12] is a four-class classification problem and the accuracy is reported as the evaluation metric. The previous methods in Table 1 include state-of-the-art end-to-end methods, including the combination of ResNet-50 with TCN [27] and the combination of EfficientNet B7 with LSTM and bidirectional LSTM [30], followed by state-of-the-art feature-based methods. The previous feature-based methods in Table 1 used different combinations of OpenFace’s eye gaze, head pose, and AU features [31] along with MARLIN’s facial embedding features [11]. These features were classified by LSTM, CNN-LSTM, TCN, and Transformer. Refer to [12] for more details. The results of the feature-based method proposed by Vedernikov et al. [34] are also reported, where the aforementioned features are classified using a Transformer-based neural network called the Tensor-Convolution and Convolution-Transformer Network (TCCT-Net).
Despite the abundant data samples in the EngageNet dataset [12] available for training complex neural networks such as ResNet + TCN [27] and EfficientNet B7 + bidirectional LSTM [30], their performance is inferior to that of feature-based methods. This highlights the necessity of extracting hand-crafted features or facial landmarks from videos and building classifiers on top of them.
@@
In the single feature configurations of previous methods in Table 1, head pose achieves better results compared to AUs, which is better than eye gaze. While head pose and eye gaze are indicators of behavioral engagement [13, 7], AUs, which are associated with facial expressions and affect [19], are indicators of affective engagement. Combining these three features is always beneficial since engagement is a multi-component variable that can be measured by affective and behavioral indicators when the only available data is video [13]. Among classification models, utilizing more advanced models improves accuracy; the Transformer is better than TCN, which is better than CNN-LSTM and LSTM; however, it comes at the cost of increased computational complexity. For single features, the Transfomer-based TCCT-Net [34] outperforms other classifiers; however, for multiple feature sets, the vanilla Transformer [12] outperforms the others.
@@
The proposed ST-GCN in Table 1, which relies solely on facial landmarks without requiring raw facial videos or multiple hand-crafted features, outperforms previous methods. Moreover, making the proposed method ordinal further improves the state-of-the-art by 3.1% compared to eye gaze, head pose, AUs, and MARLIN features [11] with the Transformer [12]. Tables 2(a) and 2(b) depict the confusion matrices of the ordinal and non-ordinal configurations of the proposed method in Table 1. As shown, incorporating ordinality significantly increases the number of correctly classified samples in the first three classes and results in a 2.7% improvement in accuracy compared to its non-ordinal counterpart.
@@
Table 2: Confusion matrices of the proposed method with (a) non-ordinal and (b) ordinal ST-GCN on the validation set of the EngageNet dataset [12].
(a) Non-ordinal ST-GCN
@@
Table 3: Classification accuracy of engagement levels on the validation set of the EngageNet dataset [12] for different variants of the proposed method.
@@
Hand landmarks are added 0.6956
The feature extraction step in [34] and [12] involves running multiple deep-learning models in OpenFace [31] to capture eye gaze, head pose, and AUs, as well as another complex network for MARLIN feature embeddings [11]. While these feature extraction networks are complex for real-time use, the proposed method relies on facial landmarks extracted using the real-time MediaPipe [14, 20]. Considering only the classification models, the number of parameters in EfficientNet B7 + LSTM [30], ResNet + TCN [27], Transformer [12], the proposed non-ordinal ST-GCN, and ordinal ST-GCN are 82,681,812, 24,639,236, 1,063,108, 861,688, and 861,431, respectively. The memory consumption for EfficientNet B7 + LSTM [30], ResNet + TCN [27], Transformer [12], and the proposed ordinal ST-GCN are 2268.78, 790.35, 178.00, and 180.96 megabytes, respectively. The inference time for classifying a data sample using EfficientNet B7 + LSTM [30], ResNet + TCN [27], Transformer [12], and the proposed ordinal ST-GCN are 348, 51, 10, and 0.8 milliseconds, respectively. This indicates the efficiency of the proposed method, which, while being lightweight and fast, also improves the state-of-the-art.
4.3.2 Variants of the Proposed Method
Table 3 displays the results of different variants of the proposed method on the validation set of the EngageNet [12] dataset. In the first two variants, the $x$, $y$, and $z$ coordinates of facial landmarks are converted into multivariate time series and analyzed by an LSTM and TCN. The LSTM includes four unidirectional layers with 256 neurons in hidden units and is followed by a $256\times 4$ fully connected layer. The parameters of the TCN are as follows: the number of layers, number of filters, kernel size, and dropout rate are 8, 64, 8, and 0.05, respectively. While their results are acceptable and better than most of the earlier methods in Table 1, they cannot outperform ST-GCN, the last two rows of Table 1. The fact that the accuracy of the LSTM and TCN in Table 3 is higher than those in Table 1 signifies the efficiency of facial landmarks for engagement measurement. In the third variant, the $z$ coordinates of facial landmarks are disregarded, and the decrease in the accuracy of the non-ordinal ST-GCN indicates the importance of the $z$ coordinates for engagement measurement. Temporal kernel sizes other than 9 in the fourth and fifth variants have a negative impact on the results of the non-ordinal ST-GCN. When engagement measurement is performed using every 2, 4, 8, and 16 frames, instead of every frame, there is a slight decrease in the accuracy of the non-ordinal ST-GCN. However, this is a trade-off between accuracy and computation since reducing the frame rate corresponds to a reduction in computation. The last row of Table 3 shows the results of the ordinal ST-GCN when 21 hand landmarks extracted using MediaPipe [14] were added to the facial landmarks. The lower accuracy compared to using only facial landmarks is due to two factors. Firstly, there is 80% missingness in the hand landmarks due to the absence of hands in the videos. Secondly, it indicates that facial landmarks alone are sufficient for engagement measurement, capturing both behavioral and emotional indicators of engagement.
4.3.3 Results on the Online SE Dataset
Table 4 presents the results of the proposed method in comparison to previous methods on the test set of the Online SE dataset [42]. The earlier methods listed in Table 4 are feature-based, extracting affective and behavioral features from video frames and performing binary classification of the features using TCN in [33], TCN in [13], LSTM with attention in [44], and BoW in [32]. Given that engagement measurement in the Online SE dataset [42] is posed as a binary classification problem, implementing the ordinal version of the proposed method is not required. The proposed method, employing facial landmarks with ST-GCN, attains the highest accuracy and AUC PR.
Table 4: Binary classification accuracy, AUC ROC, and AUC PR of engagement on the test set of the Online SE dataset [42]: comparison of previous methods against the proposed method. Bolded values denote the best results.